# Perceptions of Moderators as a Large-Scale Measure of Online Community Governance

### v1

@@ 5.4 Novice and Experienced Mods
Figure 9: In contrast to small moderator teams, large teams are much more likely to appoint new moderators who already have moderation experience. $94\%$ of mods recruited to join mod teams with more than 100 mods have at least 2 years of experience, while 74% of mods who join small teams with fewer than 10 mods have no previous experience at all.
Results. We find that large communities are much more likely to recruit moderators who already have moderation experience (Figure 9). While $74\%$ of mods appointed to communities with fewer than 10 mods are first time moderators, $94\%$ of mods appointed by communities with $>100$ mods have more than 2 years of experience.
@@
6 Discussion
The massive diversity of online communities offers enormous potential to empirically study how to make online communities better. Any such studies, however, require robust and scalable methods to quantify both the outcomes (which communities are doing ‘better’ than others), as well as the independent variables (the aspects of communities that might make them better). Measuring the success of a community’s moderators is more challenging than many other aspects of making a community ‘better,’ as unlike, say, misinformation, governance is far less visible in a community, and its success is far less well defined. The primary contribution of our work is the use community members’ own perceptions of their moderators, leveraging millions of people’s perceptions of good governance, rather than attempting to define good governance on our own terms.
We are tremendously excited about the potential synergies this line of work enables. While the findings presented in this paper are impactful by themselves, combining our measure of governance with many other important measures of community outcomes (safety, inclusion, and discussion quality, for example), along recent advances in causal inference methodology, will enable studies of how to make communities better that are both more comprehensive and more robust than previously possible.
6.1 Limitations
@@
Even though our fine-tuned sentiment classifier exceeds the performance of state-of-the-art methods (§3.3, it is not perfect. The large volume of data of data used in our analyses minimizes the impact of any single misclassification by the model, and all of our figures include bootstrapped CIs to better understand the robustness of our findings. Furthermore, systems as complex as online communities have countless confounding factors that can bias analuses such as ours. While we attempt to control for confouders by stratifying results by community size and topic and qualitatively assessing the different strata (§4), future work could make use of more sophisticated methods to control for confounders.
@@
While our methods for quantifying communities’ practices are sophisticated, they also miss many key aspects of governance. Our moderator timelines (§3.1) are limited in their temporal resolution of the end of moderators’ appointments, especially for smaller communities. Our analyses of moderators’ engagement with communities (§5.2) only consider how much each moderator posts in their communities, not what they post. Future analyses could examine the types of contributions that moderators make. Our analyses of moderator team dynamics are also limited by the lack of publicly available data about which moderators are active; many communities’ mod teams contain mods who do not contribute to the day-to-day governance of the community. Future work should incorporate detailed information about specific moderators’ actions, although data collection is a substantial challenge (Li, Hecht, and Chancellor 2022a).
6.2 Conclusion
Good governance is critical to the functioning of online communities, yet it is difficult to know what governance practices are most effective, as it is challenging to measure the ‘success’ of community governance. In this work, we developed a method to quantify community members’ perceptions of their moderators across 8,477 communities. We relate these perceptions to different aspects of governance including community size and topic (§4) as well as to different actions that moderators can take, including rule enforcement (§5.1), community engagement (§5.2), and moderator recruiting practices (§5.3-5.4). We empirically identify promising strategies for community moderators, including tailoring the strictness of rule enforcement to the community topic, and recruiting engaged mods. We make our model and datasets public to support future research.
Acknowledgements
This research was supported by the Office for Naval Research (#N00014-21-1-2154). NSF grant IIS-1901386, NSF CAREER IIS-2142794, NSF grant CNS-2025022, and the Bill & Melinda Gates Foundation (INV-004841).
@@ Appendix B Supplementary Figures
Figure 10: On average, communities with a larger fraction of their content removed by mods tend to have more a smaller fraction of their mod discourse have positive sentiment (a), and a larger fraction with negative sentiment (b). Communities with more content removed by mods also tend to have more total mod discourse (c).
@@
Figure 11: Generally, the fraction of mod discourse with negative sentiment is higher in communities with more removed content. However, these trends vary depending on the topic of the community.
@@
Figure 12: On average, adding a new moderator to a subreddit results in an increase in the fraction of mod discourse which has positive sentiment, and a corresponding decrease in the fraction that has negative sentiment. However, the magnitude of the impact varies with the size of the community’s moderator team; adding a single new mod to a community with fewer than 4 mods has a much larger impact than adding a single mod to a community with more mods. Adding a single moderator to a community with more than 15 mods has an impact which is not significantly different from 0.
Appendix C FAIR Principles and Dataset Datasheet
C.1 Compliance with FAIR Principles
Our datasets, which we will make public upon publication of this work, comply with the FAIR Principles (FORCE11 2020).
Findable. Our dataset of mod discourse makes use of existing reddit post and comment ID identifiers. We will publish them with metadata that clearly and accurately describes the data and species the data identifier on a searchable hosting platform.
Accessible. Our datasets are of small enough size ($\approx 5GB$ total) that they can be feasibly downloaded in their entirety using simple and free protocols such as https. No special authentication will be needed to access these data.
Interoperable. Our metadata will be formal, accessible, and applicable, and will use FAIR compliant vocabularies and references.
Reusable. Our datasets are widely usable, with clearly documented attributes, and will be published with an open and clearly stated data usage license.
C.2 Dataset Datasheet
Motivation. These datasets were created to understand community governance on reddit at a large scale.
Composition. Instances composing the mod discourse dataset are posts and comments discussing the moderators of the subreddit each post and comment was made in. Instances composing the moderator timelines dataset represent a moderator’s tenure in a specific subreddit. The mod discourse dataset contains 1.89 million posts and comments, while the moderator timelines dataset contains 89,237 moderator tenures. Each dataset contains all possible instances for their respective time ranges (January 2020 to June 2021 for the discourse dataset, 2006-May 10, 2022 for the timelines dataset). Each instance of the mod discourse dataset consists of a reddit-issued identifier for the associated post and comment, and a ‘predicted sentiment’ label describing the sentiment the post or comment holds towards the moderators: either positive, neutral, or negative. Instances of the mod timelines dataset consist of a moderator’s username, the subreddit they were appointed to, the exact timestamp their appointment started, and an estimated lower bound on the time their appointment ended, based on the archived snapshots the timelines were reconstructed from. To the best of our knowledge, no information is missing. The mod discourse dataset was created using a language model to classify the sentiment of posts and comments with regards to the moderators and thus may contain classification errors. Details of model performance are included in §3.3. The datasets are self contained, although they may be optionally linked to other datasets of reddit content by moderator username and/or post/comment ID. The datasets do not contain confidential or offensive information.
Collection Process. The mod discourse dataset was created by applying a labeling pipeline to raw reddit posts and comments, as described in detail in §3.3. Raw reddit data was obtained using the Pushift dataset dumps (Baumgartner et al. 2020). No sampling was performed in the production of the dataset, and no laborers were used for labeling tasks beyond the authors’ labeling of training and test sets. The timeframe of the data spans all reddit posts and comments made between December 31, 2019, and August 1, 2021. This work was approved by the University of Washington IRB under ID number STUDY00011457.
Preprocessing/Cleaning/Labeling. Several steps of preprocessing were applied to the datasets, as described in detail in §3. While due to licensing and ethical reasons we cannot republish the original raw reddit data used in their production, we make all code used for preprocessing and labeling public.
Uses. Our datasets have already been used for the analyses of reddit governance presented in this paper. A link to this paper will be clearly posted along with the datasets. These datasets can be used for many other analyses of moderation and governance on reddit. We do not believe there are any tasks for which the dataset could reasonably be used that would result in breaches of users privacy or other harms.
Distribution. Our datasets will be distributed publicly upon peer-reviewed publication of this paper. They will be published with an open and clearly stated data usage license. The datasets have no IP restrictions on them, and, as far as we are aware, are not subject to export controls or regulatory restrictions.
Maintenance. Our dataset will be hosted publicly on our University-hosted webpage, as well as on a public academic dataset hosting service to be determined at the time of publication. Contact information for the datset (the corresponding author of this paper) will be made clear and public. Our datasets are derived only from publicly-posted data which we have processed, and we will provide an affordance for users to have their data removed from our dataset.

### v2

@@ 6.4. Novice and Experienced Mods
Figure 8. In contrast to small moderator teams, large teams are much more likely to appoint new moderators who already have moderation experience. $94\%$ of mods recruited to join mod teams with more than 100 mods have at least 2 years of experience, while 74% of mods who join small teams with fewer than 10 mods have no previous experience at all.
Results. We find that large communities are much more likely to recruit mods who already have moderation experience (Figure 8). $74\%$ of mods appointed to communities with fewer than 10 mods are first time mods, while $94\%$ of mods appointed to communities with $>100$ mods have more than 2 years of experience.
@@
7. Discussion
The massive diversity of online communities offers enormous potential to empirically study how to make online communities better. Any such studies, however, require robust and scalable methods to quantify both the outcomes (which communities are doing ‘better’ than others), as well as the independent variables (the aspects of communities that might make them better). Measuring the success of a community’s moderators is more challenging than many other aspects of making a community ‘better,’ as unlike, say, misinformation, governance is far less visible in a community, and its success is far less well defined. The primary contribution of our work is the use of community members’ own perceptions of their moderators, leveraging millions of people’s perceptions of good governance, rather than attempting to define good governance ourselves.
We are tremendously excited about the potential synergies this line of work enables. While the findings presented in this paper are impactful by themselves, combining our measure of governance with many other important measures of community outcomes (safety, inclusion, and discussion quality, for example) will enable studies of how to make communities better that are both more comprehensive and more robust than previously possible.
7.1. Limitations
@@
Even though our fine-tuned sentiment classifier exceeds the performance of state-of-the-art methods (§3.3), it is not perfect. The large volume of data of data used in our analyses minimizes the impact of any single misclassification by the model, and all of our figures include bootstrapped CIs to better understand the robustness of our findings. Furthermore, systems as complex as online communities have countless confounding factors that can bias analyses such as ours. While we attempt to control for confounders by using IPTW and DID causal inference methods (§5), future work should make use of active experimentation for gold-standard causal estimates.
@@
While our methods for quantifying communities’ practices are sophisticated, they also miss many key aspects of governance. Our analyses of moderators’ engagement with communities (§6.2) only consider how much each moderator posts in their communities, not what they post. Future analyses could examine the types of contributions that moderators make. Our analyses of moderator team dynamics are also limited by the lack of publicly available data about which moderators are active; many communities’ mod teams contain mods who do not contribute to the day-to-day governance of the community. Future work should incorporate detailed information about specific moderators’ actions, although data collection is a substantial challenge (Li et al., 2022a).
7.2. Conclusion
Good governance is critical to the functioning of online communities, yet it is difficult to know what governance practices are most effective, as it is challenging to measure the ‘success’ of community governance. In this work, we developed a method to quantify community members’ perceptions of their moderators across thousands of communities. We relate these perceptions to different aspects of governance including community size and topic (§4) as well as to different actions that moderators can take, including rule enforcement (§6.1), community engagement (§6.2), and moderator recruiting practices (§6.3-6.4). We empirically identify promising strategies for community moderators, including tailoring the strictness of rule enforcement to the community topic, and recruiting engaged mods. We make our models and datasets public to support future research.
8. Ethical Considerations
We believe this work will have a positive broader impact by informing better moderation practices in online communities, as well as providing researchers with better tools to study community members’ perceptions of moderators. As we only make use of public data, we believe our work has minimal risk to participants’ privacy. As research has shown that some online community users are uncomfortable with their data being used for research, even when posted publicly (Fiesler and Proferes, 2018), we take further steps to reduce potential harms and misuse potential of our mod discourse dataset: we do not publish usernames or identifiable information, only predicted sentiment with regards to the moderators. We publish moderator timelines, including moderators’ usernames, however these usernames are already publicly listed on communities’ ‘about’ pages. Upon publication of this dataset, we will provide affordances for users to have their data removed at their request. This study was approved by the University of Washington IRB under ID number STUDY00011457.
Supplementary Materials
@@ Appendix B Supplementary Figures
Figure 9. On average, communities with a larger fraction of their content removed by mods tend to have more a smaller fraction of their mod discourse have positive sentiment (b), and a larger fraction with negative sentiment (c). Communities with more content removed by mods also tend to have more total mod discourse (a).
@@
Figure 10. Generally, the fraction of mod discourse with negative sentiment is higher in communities with more removed content. However, these trends vary depending on the topic of the community.
@@
Figure 11. On average, adding a new moderator to a subreddit results in an increase in the fraction of mod discourse which has positive sentiment, and a corresponding decrease in the fraction that has negative sentiment. However, the magnitude of the impact varies with the size of the community’s moderator team; adding a single new mod to a community with fewer than 4 mods has a much larger impact than adding a single mod to a community with more mods. Adding a single moderator to a community with more than 15 mods has an impact which is not significantly different from 0.
Appendix C Covariates Used in Propensity Score Modeling for IPTW
C.1. Covariates used for Mod Workload Analyses
The following table shows the covariates that were used in a logistic regression propensity score model for the moderator workload analyses (Figure 4), along with the resulting Standardized Mean Differences (SMDs) after reweighting, and the associated propensity score model (P.S.) coefficients used in the logistic regression.
Moderator Workload Treatment Bin (Posts+Comments per Mod per Day)
0-5 5-10 10-100 ¿100
SMD P.S. Coefficient SMD P.S. Coefficient SMD P.S. Coefficient SMD P.S. Coefficient
Covariate
total_items -0.18 -0.55 -0.17 -0.45 -0.10 +0.07 +0.48 +0.93
frac_deleted +0.05 +0.01 +0.02 +0.02 -0.07 -0.07 +0.20 +0.04
frac_removed +0.19 +0.04 +0.07 +0.01 -0.06 -0.05 -0.03 +0.00
num_mods +0.04 +0.29 -0.02 +0.23 -0.03 -0.08 -0.01 -0.44
category_hobby -0.22 -0.02 +0.01 +0.01 +0.06 +0.02 -0.07 -0.01
category_discussion +0.18 +0.01 -0.02 -0.01 -0.05 -0.01 +0.04 +0.01
category_memes +0.18 +0.03 +0.01 +0.00 -0.04 -0.05 +0.00 +0.01
category_news +0.03 -0.01 +0.01 +0.00 -0.01 +0.01 +0.00 -0.01
category_media +0.34 +0.02 +0.10 +0.00 -0.08 -0.04 -0.00 +0.01
category_identity -0.24 -0.04 -0.10 -0.01 +0.05 +0.07 +0.07 -0.02
C.2. Covariates used for Strictness of Rule Enforcement Analyses
The following table shows the covariates that were used in a logistic regression propensity score model for the rule enforcement analyses (Figure 5), along with the resulting Standardized Mean Differences (SMDs) after reweighting, and the associated propensity score model (P.S.) coefficients used in the logistic regression.
Amount of Removed Content Treatment Bin (Percentage of All Content)
0%-1% 1%-2% 2%-3% ¿3%
SMD P.S. Coefficient SMD P.S. Coefficient SMD P.S. Coefficient SMD P.S. Coefficient
Covariate
total_items -0.10 -0.12 +0.03 -0.10 +0.12 +0.01 +0.12 +0.21
frac_deleted -0.58 -0.51 +0.02 -0.01 +0.30 +0.10 +0.67 +0.42
mod_workload -0.08 -0.00 +0.04 +0.14 +0.15 +0.02 +0.04 -0.16
num_mods -0.06 -0.21 -0.01 +0.12 -0.01 +0.01 +0.16 +0.08
category_hobby +0.53 +0.10 -0.17 -0.01 -0.29 -0.01 -0.46 -0.08
category_discussion -0.26 -0.09 +0.00 +0.01 +0.02 -0.01 +0.40 +0.09
category_memes -0.04 +0.08 -0.01 -0.00 +0.04 +0.01 -0.02 -0.08
category_news -0.11 -0.12 -0.06 -0.06 -0.00 -0.00 +0.28 +0.19
category_media -0.12 +0.03 -0.08 -0.01 +0.02 -0.00 +0.21 -0.02
category_identity -0.33 +0.00 +0.32 +0.08 +0.31 +0.02 +0.02 -0.10